Father age now 32 yrs, son age is 7 yrs, in how many yrs father will be twice of son age.
Answer by sachi(548)

Father age now 32 yrs, son age is 7 yrs,
say in x yrs father will be twice of son age.
so 32+x=2(7+x)=14+2x
or 2x-x=32-14
or x=18
so after 18 yrs father will be twice of son's age